How many grams of methane gas (CH4) are in a 15.8-liter sample at 1.2 atmospheres and 27°C? Show all work used to solve this problem.
1
Expert's answer
2018-07-11T14:40:08-0400
T=27℃=300.15K.
R=0.082057 ((l∙atm))⁄((mol∙K)),
Due to Ideal gas law: PV=nRT,
n=PV/RT=(1.2atm×15.8l)/(0.082057 ((l∙atm))⁄((mol∙K))×300.15K)=0.7698moles,
So,n(CH_4 )=0.7698moles,
m(CH_4 )=(1.008×4+12.011) g⁄moles×0.7698moles=12.35g.
Answer: m(CH_4 )=12.35g.